OpenStudy (anonymous):

A line with a slope of -2 crosses the y-axis at (0, 3). The equation of the line is

OpenStudy (anonymous):

its the same as before

OpenStudy (anonymous):

y=mx+b

Nnesha (nnesha):

y = mx +b again slope is m y -intercept point is where line cross y-axis when 

OpenStudy (anonymous):

m is the slope plug it in

OpenStudy (anonymous):

y=3x+b

OpenStudy (anonymous):

y=-2x+b

Nnesha (nnesha):

yes you can substitute values or just use definition of y-intercept

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

3=-2(0)+b 3=B...

Nnesha (nnesha):

yes you can substitute values or just use definition of y-intercept y-intercept where line cross y-axis when x =0

OpenStudy (anonymous):

0+b

Nnesha (nnesha):

yes that's right so when x =0 y = y-intercept

Nnesha (nnesha):

so when you look at the order pair if first number is 0 2nd number of order pair would be y=intercept always ! :-)

OpenStudy (anonymous):

I don't get it..

OpenStudy (anonymous):

I get that the y intercept it -2

OpenStudy (anonymous):

I mean 3

OpenStudy (anonymous):

And the slope is -2

Nnesha (nnesha):

okay  y-intercept is a point where line cross y-axis